A potential difference of 100 V is applied to the primary coil of a transformer. The current through the primary coil is 2.0 A. The potential difference across the secondary coil is 15 V


Calculate the current flowing through the secondary coil.


1
Expert's answer
2022-03-25T09:11:19-0400

Let us assume transformer is ideal, so

"P_s=P_p\\
I_sV_s=I_pV_p"

The current in the secondary coil

"I_s=\frac{I_pV_p}{V_s}=\frac{2.0\times 100}{15}=13\:\rm A"